OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (e.mccormick):

Do you know the ratios for a 30-60-90 traingle? http://mathworld.wolfram.com/30-60-90Triangle.html

OpenStudy (e.mccormick):

Because the ratios are set, you can find all sides from any one side.

OpenStudy (anonymous):

I'm sorry, I'm still confused on how to find the answers @e.mccormick

OpenStudy (e.mccormick):

OK. They give you the side oppoite the 30, right?

OpenStudy (e.mccormick):

If we do the ratop as 30:60:90 as in side opposite, then the diagram ratio becomes: \(\dfrac{a}{2}:\dfrac{a\sqrt{3}}{2}:a\) You are given: \(41:x:y\) Now, because you know the first one, it is best... easiest if the ratio you work with has 1 in that spot. So I am gong to multiply through the first ratio with 2. \(a:a\sqrt{3}:2a\) Ah ha! Now if you use 41 as a, what do the other two values become?

OpenStudy (anonymous):

I'm still quite confused. Sorry math isn't my strong suit... @e.mccormick

OpenStudy (e.mccormick):

Replace a with 41 and multiply.

OpenStudy (e.mccormick):